From GlueXWiki
Jump to: navigation, search

February 25, 2016 Drift Chamber meeting


  1. Instructions for Bluejeans meeting connection
  2. Meeting ID: 290664653
  3. To join via a Web Browser, go to the page [1] https://bluejeans.com/290664653.


  1. NIM Papers
  2. Run Status Run Coordination Meetings: Spring 2016 Run
  3. Electronics (Fernando, Chris, Nick)
  4. Other


Participants: Curtis, Naomi, Mike (CMU) Luke, Beni, Chris, Cody, Simon, Nick, Sergey, and Lubomir (JLab).

NIM papers

- As Naomi proposed for the CDC paper: Beni will write the Introduction and the Conclusions. Naomi, Mike,Simon will work out everything in between. Curtis will proofread the paper. Naomi wants to use the "old" svn system working on one file (section) at a time without interfering with somebody else.

- Lubomir is working on the VCI contribution (4 pages in NIM) to be submitted by middle of March.

Run Status

- Beni: the HV channel in CDC that was problematic is now OK. The corresponding cards are oriented horizontally and are also at the top side, so most likely some kind of "dust" falls on them causing this problem.

- We see a factor of 4-5 higher currents now (both in the FDC and the CDC) compared to spring 2015 - see log entry linked above. We had half the energy at that time, but that's not likely enough to explain the difference. To understand the origin of the background we want to have measurements with empty target. We had it once but there are problems with the archiver for that period: the target status had wrong timing stamp.

- Two FDC HV channels (same cell4- hv4 for both, package 2 and 3) show strange behavior: they are OK as long as there's no beam, but with beam we see current jumps (plot attached) up to 5 uA for both positive and negative HV. This indicates that something is going on inside the gas volume, somehow triggered by the beam.

- Lubomir studied the average amplitude as a function of the drift time for two run periods - fall 2014 and spring 2015 (plot linked above). There's a drop of 30% for both periods. The current data is being analyzed. The plan is also to look at the data from the full test of the third package in the EEL126, but it will take some time to adjust the software for that. Mike will do the same analysis for the CDC.

fADC125 firmware

Naomi added this:

Cody put some more timing constraints into the firmware and and compiled it into a new version 2.09. He and Beni tested it in the hall on Tuesday and Naomi revived the prototype setup at CMU and tested it there, this works but it is slow (only 24 channels, and slow data transfer off the computer). Beni and Cody found that the duplicate sample words in the long readout mode are still present. However, most of the errors in the calculated quantities due to timing problems have been corrected. Naomi found 0 errors in leading edge times from 547060 hits. There are a few logic issues in the peak search and possibly also a timing issue. We will discuss this further and try again. One of the peak search issues occurred when there were several samples of value 0 before the peak, because Naomi did not set the pedestals well.